OpenStudy (anonymous):

Polynomial help? (4x2 + 15x – 3) – (–3x2 + 5)

OpenStudy (zpupster):

change signs inside brackets of second parentheses 3x^2 - 5 and then remove all brackets combine like terms 4x^2+ 3x^2= 15x= and -3-5=

OpenStudy (anonymous):

My choices are: A. 7x^2 + 15x – 8 B. x^2 + 12x + 2 C. x^2 + 15x – 8 D. 7x^2 + 8

OpenStudy (anonymous):

So how do I get your equation to fit into one of those choices?

OpenStudy (zpupster):

4x^2+3x^2 = 7x^2 15x=15x and -3-5=-8 write these out like so 7x^2+15x-8

OpenStudy (anonymous):

Oh. Thanks!